OpenStudy (anonymous):

For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p? A. (x, x + 4) B. (x, 3p + 4) C. (3p, x + 4) D. (3p, 3p + 4)

OpenStudy (anonymous):

@Mesopretty

OpenStudy (anonymous):

Need help huh?

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

f(3p) = x + 4 f(3p) = 3p + 4 (3p, 3p+4) x = 3p y= 3p + 4

OpenStudy (anonymous):

so c ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thank you :*

OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 19 16 8

OpenStudy (anonymous):

Substitute the values. 3. g(x) = x^2 + 3, find g(4) So, 4 = 4^2 + 3. Can you simplify from there? 4. If h(x) = 2x + 1, find the value of x if h(x) = 11 So, 11 = 2x + 1 and solve for x.

OpenStudy (anonymous):

Its alright, just solve the equation... For #2 you need to find the value of x. 11 = 2x + 1 Subtract 1 from both sides. 11 - 1 = 2x 12 = 2x divide by 2. x = 6